OpenStudy (katecc379):

Find the diameter of a cone that has a volume of 104.67 cubic inches and a height of 4 inches. Use 3.14 for pi.

OpenStudy (katecc379):

@TheSmartOne

OpenStudy (david-elijah):

im tired of you

OpenStudy (thejoker):

what is the radius?

OpenStudy (katecc379):

does not give it

OpenStudy (thejoker):

Surface Area?

OpenStudy (katecc379):

it only gives whats in the question i put

OpenStudy (thejoker):

I am getting approx. 78.48079589569615

OpenStudy (katecc379):

thats none of the answers

OpenStudy (katecc379):

4 inches 6 inches 8 inches 10 inches

TheSmartOne (thesmartone):

volume of a cone = 1/3 * height * radius^2 * pi

TheSmartOne (thesmartone):

to find the radius radius * 2 = diameter

OpenStudy (katecc379):

still confused

TheSmartOne (thesmartone):

ok, can you list what you've been given? and what you need to solve for?

OpenStudy (katecc379):

volume= 104.67 height=4in solve for diameter

TheSmartOne (thesmartone):

ok, now let's plug that in to our formula first we'll need to solve for radius, and then we can find the diameter by simply multiplying the radius by 2 104.67 = 1/3 * 4 * radius^2 * 3.14 can you simplify it? :P

OpenStudy (katecc379):

i got 5

OpenStudy (katecc379):

@TheSmartOne

TheSmartOne (thesmartone):

104.67 = 1/3 * 4 * radius^2 * 3.14 (104.67 * 3)/(4 * 3.14) = r^2 r^2 = 25.0007962 r is approximately 5 great job! But we need to find the diameter, so remember this: diameter = 2 * radius plug in the radius and get the diameter :)

OpenStudy (katecc379):

so 10 could you help with one more?

OpenStudy (katecc379):

@TheSmartOne

TheSmartOne (thesmartone):

mhmm, http://www.mathopenref.com/constparallel.html an interactive thing that shows you how to construct a parallel line with a compass what do you think the next step will be? :)

OpenStudy (katecc379):

A?

TheSmartOne (thesmartone):

correct :)